OpenStudy (anonymous):

Hey guys, anyone know the answer to these? What is the solution of the equation? 2.8 = 2y A. 1/4 B. 1/2 C. 2.8 D. 0.6 What is the solution of the equation? -9 = 5/7n A. -153/5 B. -5/153 C. -17/45 D. -45/17

OpenStudy (anonymous):

The solution is none of those, the solution for y would be 1.4

OpenStudy (anonymous):

This is incorrect, there is no right answer. You would divide both sides by 2 2.8=2y 2.8/2=y 1.4=y The second one is also incorrect. -9 = 5/7n -9/(5/7)=n -12.6 or -63/5=n

OpenStudy (anonymous):

When you have a number on one side and a veritable (letter) and a number on another, you always want to get the letter by itself. You do this by dividing the number on both sides.

OpenStudy (anonymous):

Okay, I get what your saying, Thank you:)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

2.8 = 2y y = \(\frac{ 7 }{ 5 } = 1.4\)

OpenStudy (anonymous):

Thank you some_someone and alexisjm1, y'all were helpful too:)